Effie and kirsten live 23.6 km apart. They decided to cycle to the pool at the park, which is located between their homes. If effie lives 5.2 km closer to the park, how far did they each cycle?
Answer by ankor@dixie-net.com(22740)

Effie and kirsten live 23.6 km apart.
They decided to cycle to the pool at the park, which is located between their homes.
If effie lives 5.2 km closer to the park, how far did they each cycle?
:
let d = dist that k is from the pool
e is 5.2 km closer, therefore
(d-5.2) = dist that e is from the pool
:
the two distances = 23.6, the dist they are from each other
therefore
d + (d-5.2) = 23.6
2d = 23.6 + 5.2
2d = 28.8
d = 28.8/2
d = 14.4km is k's distance from the pool
and
14.4 - 5.2 = 9.2 km is e's distance
:
:
check: 14.4 + 9.2 = 23.6